OpenStudy (anonymous):

A=a+b+c / over 3 solve for b A.) 3A - a - c = b B.) -c / 3 = b C.) A - c / 3 = b D.) -3 - c = b

OpenStudy (anonymous):

First get rid of the denominator by multiplying both side by 3. Then solve for b by moving a and c to the other side leaving b alone.

OpenStudy (anonymous):

A(3) = (a+b+c)/3 * 3 3A = a+b+c 3A-a-c = a+b+c-a-c 3A-a-c = b b= 3A-a-c